# Unbundle-Rewrite-Rebundle: Runtime Detection and Rewriting of Privacy-Harming Code in JavaScript Bundles

### v2

@@ 4.4. Performance
Considering URR’s ability to recognize bundles given the entire contents of scripts, we determined that an approach that intercepts network requests and responses will be a useful initial deployment. Considering Chrome’s restrictions on content filtering from extensions (Dunk, 2023), we developed a Firefox extension that intercepts and modifies the content of responses across all domains that the browser visits (Wu, 2022).
@@
    Next, we used Mozilla’s web-ext (Web, 2023) to load a Firefox instance with the web extension.
@@
    We used puppeteer (Pup, 2023) to connect to the browser instance, and visit the web page in a new tab and wait for 30 seconds.
@@
    The extension’s background script intercepts and evaluates all script-based network requests. It captures and stores the time taken by each step in the script evaluation pipeline (MDN, 2023a).
@@
Table 3. Time taken (ms) by the the extension to process scripts using network interception.
@@
First, we discuss the evaluation of individual scripts. Table 3 shows the time taken by different phases of the pipeline. We observed the gathering modules and comparing processed modules have the least impact, taking on average 46.17ms and 1.15ms respectively. URR takes the longest time to process modules, i.e., recursively compute hashes for all nodes in an AST, which takes 326.31ms on average for each script.
However, we observed that the largest chunk of the time was consumed outside of URR’s evaluation and instead spent in buffering incoming response chunks to load the entire script. This is a limitation of the mode of deployment, i.e., Firefox extensions that modify the body of network responses bypass the browser’s optimized cache for scripts (MDN, 2023b). The deployment waits for large network responses to complete before evaluating the script. In Section 5, we discuss alternative deployments that evaluate code at runtime and can bypass dependence on network load times.
@@
First Contentful Paint (FCP) is a timing measure that shows when the browser renders the first bit of content. The content could be any text, image, video, canvas, or non-empty SVG. The timing shows the first instance that the user has an indication that the page is loading. We observed that the extension added 293.36ms to the average time a user would wait for such an indication.
@@
DOM Interactive indicates the time taken for the Document Object Model (DOM) parser to finish its work on the main document, i.e., the time taken to construct the DOM. This time can be affected by the parser-blocking JavaScript. Note that the extension’s script evaluation runs outside the main thread. We observed that the extension added 571.66ms on average, i.e., a user would have to wait an additinal 0.5s before the browser has parsed the DOM.
@@
Page Load indicates the total time taken for all resources to load. It indicates that network requests for all scripts, images, and other resources have completed. This metric does not indicate actual end-user experience since it depends on device capabilities and various network conditions, but in our case, provides insight into the additional time added by the extension buffering responses and evaluating script contents. We observed that pages loaded with the extension took an additional 2,317.06ms on average before triggering the load event.
Overall, deploying URR as a browser extension slightly increases the time taken to load the page, adding minimal performance overhead and ensuring that users can continue to meaningfully interact with the website.
@@ 5. Discussion
5.1. Applicability to Other Systems
@@
In this work we target and evaluate against JavaScript bundles generated by Webpack. We selected Webpack because it is the most popular bundling system on the Web. However, there are many other bundling approaches used on the Web, some variations on the same approached used in Webpack some fundamentally different and even working on different levels of the deployment process. We here briefly discuss these other bundling formats, and how URR could be extended to apply to them (and with what difficulties).
@@
Webpack is generally used to combine multiple different JavaScript libraries and code units, and to process them into a single JavaScript file, to make deployment easier and (in some cases) execution faster. While Webpack is the most common tool for this purpose, there are many others, for use with different languages, build chains, testing frameworks, and, in some cases, to also bundle additional resource types beyond JavaScript files. Examples of these alternative bundlers include Browserify  [^1] https://browserify.org/ and Gulp  [^2] https://gulpjs.com/, among many others.
@@
URR could easily be extended to cover these other bundling tools, as at root they all operate in the same manner (i.e., consume multiple JavaScript files, preprocess them, and then generate code for the resulting combined AST). To do so would only require generating new signatures for ASTs of each target privacy-harming library generated by the bundler’s preprocessing and rewriting phases, and understanding the structure of each bundler combines the ASTs generated by each input library into the final, resulting code unit. This is work that would only need to be done once per bundler version, and then could be shared across all URR clients.
@@
Another approach to JavaScript bundling is to directly combine each JavaScript code unit into a single archive, and to ship the entire archive to the client, along side the website’s initial HTML. This approach, exemplified by Google’s WebBundles  [^3] https://wpack-wg.github.io/bundled-responses/draft-ietf-wpack-bundled-responses.html proposal, does not preprocess or otherwise modify the include JavaScript code units; included files are directly copied into the bundle archive. Cloudflare’s Cloudflare’s Managed Components  [^4] https://managedcomponents.dev/ product can also be seen as a form of this kind of bundling, though instead of all combined into a single archive, they’re instead delivered “on demand”, as managed by an overriding “manager” application, making any URLs unpredictable.
Extending URR to cover these kinds of bundled application would be trivial, since bundled JavaScript files are not modified or preprocessed in the bundling process. Identifying them within the bundled application is straightforward. Similarly, rewriting unwanted code is similarly trivial, since it only requires swapping the original file with the privacy-preserving alternative (either in the original archive, or returned as a new subresource by the “manager” application).
5.2. Alternative Deployment Strategies
@@
In this work we implement URR as a Firefox extension. We choose a Firefox extension because only Firefox’s extension API includes the ability for an extension to buffer and rewrite a fetched subresource  [^5] https://developer.mozilla.org/en-US/docs/Mozilla/Add-ons/WebExtensions/API/webRequest/filterResponseData (like a JavaScript file), before the file is seen, parsed, and executed by the JavaScript engine.
@@
However the same approach we take in the extension could be identically deployed from other decision points, either to be more general across browsers (e.g., as part of a man-in-the-middle proxy) or more specific to particular browsers (e.g., as a modification to the V8 parsing pipeline). The majority of the resources needed for URR to work could apply equallyacross all of these intervention points (e.g., fingerprints of the ASTs of unwanted to code, bundle “debundling” logic, privacy-preserving replacement AST subtrees).
@@
However, each of these intervention points would come with their own largely-predictable trade offs. Analyzing, debundling, and rewriting JavaScript bundles as part of a man-in-the-middle proxy would work for any browser or network tool, though with a performance impact (since JavaScript engine optimizations like partial or lazy compilation would not be possible). Likewise, pushing URR logic into the JavaScript engine directly would likely allow for greater performance, at the cost development and maintenance cost, and requiring browser specific implementations. That said, we reemphasize that most of the novel aspects of URR would be generic and shareable across all possible intervention points.
5.3. Diversity of Target Library Representation
@@
URR requires a significant amount of precomputation to work effectively. Specifically, URR requires precomputing a signature for every AST for each library or code unit that should be rewritten at runtime. While this is a significant improvement over the existing state of the art ((Smith et al., 2021), for example, requires the precomputing each target bundled application, which will both be orders of magnitude larger in occurrence, but also enormously larger in terms of required disk space), its still not trivial. The same target library can give different signatures depending on library version (e.g., FingerprintJS v2 vs v3), bundler version (e.g., Webpack v4 vs Webpack v3), bundler optimization strategies (e.g., “tree-shaking” or no), library integration method (e.g.,CommonJS vs ECMAScript modules) among other dimensions.
@@
As discussed in Section 3, URR uses several heuristics to effectively generalize signatures, to flatten the number of dimensions of signatures needed per target library (e.g., label stripping, AST simplification, etc.). Never the less, there is still a tradeoff between coverage—generating as many signatures as possible, to correctly identify the same target code across a wide range of representations—and concision—minimizing the memory, matching time, and disk space used on each URR client at runtime.
5.4. Replacements
@@
While other phases of URR are programmatically generated, we manually created benign replacements. This part requires an understanding of both, the specific bundler and the target library itself. While prior work has shown approaches that can automate this creation, we leave the adaption of a similar approach in the context of bundles as an avenue to explore in future work. Prior work like (Smith et al., 2021), which allows for the automatic creation of privacy-and-compatibility preserving versions of JavaScript libraries, could be leveraged to greatly expand the number of target libraries URR can identify and rewrite in bundled applications.

### v3

@@ 4.3. Performance
Considering URR’s ability to recognize bundles given the entire contents of scripts, we determined that an approach that intercepts network requests and responses will be a useful initial deployment. Considering Chrome’s restrictions on content filtering from extensions [^30] https://developer.chrome.com/blog/improvements-to-content-filtering-in-manifest-v3/, we developed a Firefox extension that intercepts and modifies the content of responses across all domains that the browser visits [^31] https://blog.mozilla.org/addons/2022/05/18/manifest-v3-in-firefox-recap-next-steps/.
@@
    Next, we used Mozilla’s web-ext [^32] https://extensionworkshop.com/documentation/develop/getting-started-with-web-ext/ to load a Firefox instance with the web extension.
@@
    We used Puppeteer [^33] https://pptr.dev/ to connect to the browser instance, visit the web page in a new tab and wait for 30 seconds.
@@
    The extension’s background script intercepts and evaluates all script-based network requests. It captures and stores the time taken by each step in the script evaluation pipeline.
@@
Table 3. Time taken (ms) by the extension to process scripts using network interception.
@@
First, we discuss the evaluation of individual scripts. Table 3 shows the time taken by different phases of the pipeline. We observed the gathering modules and compared processed modules, which have the least impact, taking, on average, 46.17ms and 1.15ms, respectively. URR takes the longest time to process modules, i.e., recursively compute hashes for all nodes in an AST, which takes 326.31ms on average for each script.
However, we observed that the largest chunk of the time was consumed outside of URR’s evaluation and instead spent in buffering incoming response chunks to load the entire script. This is a limitation of the mode of deployment, i.e., Firefox extensions that modify the body of network responses bypass the browser’s optimized cache for scripts [^34] https://developer.mozilla.org/en-US/docs/Mozilla/Add-ons/WebExtensions/API/webRequest/filterResponseData. The deployment waits for large network responses to complete before evaluating the script. In Section 5, we discuss alternative deployments that evaluate code at runtime and can bypass dependence on network load times.
@@
Table 5. Memory space taken by the extension’s hash signatures to detect target libraries, comprising signatures of target libraries along with those of additional libraries (§4.1.4).
Signatures #ASTs #HashMap Size
Target Addtional Entries
All Hashes All Hashes 258,788 265,962 3.55 GB
All Hashes 0.5 * (All Hashes) 176,963 142,173 2.77 GB
All Hashes 95,174 19,555 2.27 GB
Top-10K Matches 242 2,514 0.003 GB
@@
First Contentful Paint (FCP) is a timing measure that shows when the browser renders the first bit of content. The content could be any text, image, video, canvas, or non-empty SVG. The timing shows the first instance in which the user has an indication that the page is loading. We observed that the extension added 293.36ms to the average time a user would wait for such an indication.
@@
DOM Interactive indicates the time taken for the Document Object Model (DOM) parser to finish its work on the main document, i.e., the time taken to construct the DOM. This time can be affected by the parser-blocking JavaScript. Note that the extension’s script evaluation runs outside the main thread. We observed that the extension added 571.66ms on average, i.e., a user would have to wait an additional 0.5s before the browser has parsed the DOM.
@@
Page Load indicates the total time taken for all resources to load. It indicates that network requests for all scripts, images, and other resources have been completed. This metric does not indicate actual end-user experience since it depends on device capabilities and various network conditions, but in our case, provides insight into the additional time added by the extension buffering responses and evaluating script contents. We observed that pages loaded with the extension took an additional 2,317.06ms on average before triggering the load event.
4.3.5. Signature Database Size.
URR’s extension deployment uses a complete database of hash signatures, which includes those computed for target library configurations (based on Table 2) along with additional configurations of popular NPM libraries (see §4.1.4). To optimize the time taken to compare modules (§3.3), URR stores the signatures as a hashmap, with the key for each entry holding the hash of a top-level node and the value holding a dictionary comprising root node hashes along with corresponding weights (see Algorithm 2). We evaluated different database compositions to reduce its overall size, especially considering subsets of additional configs from popular libraries. Table 5 shows that even with the database solely comprising signatures of target libraries, it consumes 2.27GB of memory space (see Row 3 of Table 5). We observed that reducing the number of hashes included in the database reduced the extension’s use of memory space (from 3.55 GB to 2.27 GB), which can be beneficial to devices with limited resources.
We considered further reducing the database to only include target library signatures we found in our 10K crawl (see §4.2.5). While this subset greatly reduced the size of the database (3.54 MB), we noted that the majority of the ASTs (57%; $n=139$) correspond to a single match. We comprised an extensive list of hash representations (based on Table 2) to make URR especially effective in detecting target libraries despite the websites adopting of a wide array of potential configurations. Using a match-based subset (as used in Row 4 of Table 5) will limit URR’s detection of libraries on websites that deploy target libraries within differently configured webpack bundles. We note that while greatly reducing target library hash representations may not be effective, developers can consider caching frequently encountered bundles. We further discuss caching in §5.
Finally, we repeated crawls of sites from the Tranco list (see §4.3.2), but this time, visited sites with each of the four database sizes deployed within URR. We observed that the different configurations had a negligible effect on page load, with the databases’ use in the Comparing Modules phase taking an average time of 1ms ($M=0$ms). Table 5 therefore focuses on the different configurations’ impact on memory usage.
4.4. Replacements
Prior work has extensively studied and determined the creation and use of non-breaking, benign replacements for privacy-harming code (Smith et al., 2021). We follow similar principles and extend example replacements used by uBlock Origin [^35] https://github.com/gorhill/uBlock/tree/9123563895f0499849b4d85c4f95e1ed6ace2231/src/web_accessible_resources.
4.4.1. Creating Benign Versions of Target Libraries.
For each of the three libraries, we considered associated source code and documentation to create equivalent benign versions.
FingerprintJS (v3 and v4), when loaded as a library from its npm package, provides a function, load(), which returns a Promise that resolves to an object that can be used to get() a unique visitorId for the user. While the computation includes privacy-harming code, the benign replacement that we developed only returns a randomly generated visitorId. We used a similar approach to the existing replacement included within uBlock origin [^36] https://github.com/gorhill/uBlock/blob/9123563895f0499849b4d85c4f95e1ed6ace2231/src/web_accessible_resources/fingerprint3.js, but modified the same to a Node.js module. FingerprintJS v2 has developers get() a unique visitorId without first needing to load() the library. Our benign replacement exports values for all versions of FingerprintJS without needing to individually support 52 minor versions released across 3 major versions (v2, v3, and v4). Considering the small number of export replacements, manually creating a replacement for FingerprintJS only takes around 20 minutes.
Sentry provides a browser library (@sentry/browser). Below, we detail the development of replacements for the library.
  1. (1)
    Enumerating Replacements. Sentry’s browser library further imports related packages, which include utilities for tracing and profiling user action events and functionality to handle network requests, along with 24 functions (with void return values) from its core analytics library (@sentry/core). We used the official documentation as a guideline. Enumerating all imports, functions, and objects that needed replacements took us around 90 minutes of wall clock time.
  2. (2)
    Functions. We used a benign function that performed no operations (noop) as a stand-in for functions with no return values. We used the functions for the 24 functions imported from the core library, 3 functions imported to handle network requests, and one function each to support its additional “hub” extension and to initialize the library. The library additionally exposes a function, close() (used to flush pending events), that returns a Promise that resolves with a timeout. We retained this function as in the original library. Once listed, since all functions required straightforward replacements, this took around 20 minutes.
  3. (3)
    Class Definition. Sentry defines a BrowserClient class with a constructor that accepts any argument it is passed. Developing a replacement for this class only takes a couple of minutes.
  4. (4)
    Global Variable. Finally, Sentry exposes all its functionality by defining an object instance of its BrowserClient class and exposes all aforementioned functions as members of this class. The library makes the object available as a global variable (on the window object). Also, it exports the same functionality to the user’s code that imports it as an npm package. Defining the global variable and the exports takes under 5 minutes.
Updates between Sentry versions (v5-v8) reuse exports of the same name, ensuring that, unlike for FingerprintJS, our benign replacements did not need to accommodate additional changes. As a result, the benign replacement that we developed in around 2 hours by wall clock time replaces 335 minor versions of Sentry’s npm package.
[…]
While Prebid’s library itself has undergone numerous changes across versions (e.g.,support for adapters, consent management, and user ID modules), its publisher-facing API has retained the same exports. We verified this against the library’s release notes [^38] https://docs.prebid.org/dev-docs/pb9-notes.html and the library’s source code. As a result, developing benign replacements for 41 functions and 4 arrays took only about an hour by wall clock time. Crucially, our replacement works against 186 minor library versions across 4 major Prebid version releases (v5-v9).
4.4.2. Limiting Breakage.
We designed our replacements in a manner that would cause little breakage. Our replacements are expertly developed to maintain compatibility (uBO uses a similar approach for non-bundle replacements, e.g.FPJS [^39] https://github.com/gorhill/uBlock/blob/9123563895f0499849b4d85c4f95e1ed6ace2231/src/web_accessible_resources/fingerprint3.js). For each replacement module, we created a barebones application that accesses the library functionality. The barebones application helps evaluate each replacement, ensuring that the application remains functional. We package the resulting application and gather the webpack module corresponding to the benign replacement for use in  the system’s deployment. We evaluated our replacements by (i) ensuring that the JS execution of mock apps that bundle the library and the replacement were both successfully executed, (ii) our deployment returned metrics showing successful page load and DOM interaction events (Smith et al. (Smith et al., 2022) found network behavior highly predictive of breakage. Nisenoff et al. (Nisenoff et al., 2023) found users consider page load and responsiveness the most prominent breakage category.) We ensured our replacements have the shape (scope and exports) we want and are effective (execution of mock apps, page load, DOM responsiveness in the real world) against extensive testing. We can add a manual breakage evaluation by sampling modified websites.
4.4.3. Manual Breakage Analysis.
Core to understanding whether a website is functional or broken is the user’s subjective perspective of the features that are needed within the context of the website. With this in mind, we adopted the methodology presented by Snyder et al. (Snyder et al., 2017) to study the effect of URR’s replacements on website functionality. We sampled 30 websites (10 sites per target library) that bundled target libraries and on which are extension deployed benign replacements (determined from the crawl in §4.2.5). We instructed three undergraduate workers to visit each website twice in a row. In the first visit, they were instructed to use the website in an unmodified Firefox browser. Considering that the average dwell time for a user on a website is under a minute (Chao et al. (Liu et al., 2010)), the workers were instructed to perform as many actions on the page as possible in one minute. On the second visit, they were instructed to visit and use the same website within a Firefox browser with the extension enabled. We asked them to assign a score for each website: 1 if there was no perceptible difference between the two visits, 2 if the browsing experience was altered during the second visit, but they were still able to complete the same tasks as during the first visit, and 3 if they were unable to complete the same tasks as during the first visit.
We considered a website to be broken if the workers could not complete their intended task (i.e., the website was assigned a score of 3). We had the workers browse the same sites and record their scores independently and also had them include notes to provide context for their scores. We observed a high agreement ratio (96.67%) between the workers’ scores, who reported that 29 of the 30 websites were functional with benign replacements. For the single site that the workers reported broken (https://mlb.com), we visited the site with a modified version of our extension that does not perform replacement. We observed that the site was broken even without the replacements, primarily owing to longer wait times in loading large scripts as fragments and the extensions’ blocking network requests (see §4.3.3). Our observations were also in line with the workers’ notes, which indicated that they considered the page broken because it takes “too long to load”. We, therefore, concluded that the site was broken due to the extension’s network interception and not the benign replacements.
We note that we did not collect or process private data or identifiable information from our evaluators, and the testing does not qualify as human subject research. Therefore, we did not seek an IRB review. This is consistent with prior work that used roughly the same evaluation setup (Iqbal et al., 2020; Snyder et al., 2017; Iqbal et al., 2021).
4.4.4. Automated Breakage Analysis.
We also performed an automated analysis by crawling websites from the Tranco Top-1K. We visited each site within a Firefox browser with and without deploying URR. We used a puppeteer script to capture all console errors and page errors that were observed during each visit. Visits with URR deployed observed an average of 5 errors ($\sigma=15$;$M=1$), while visits without URR observed a slightly higher but similar average of 5.66 errors ($\sigma=14$;$M=1$). We noted that visits to 84% sites ($n=$714/846) with URR observed the same number of errors or fewer. Interestingly, we found more sites that produced fewer errors with deploying URR ($n=$152;18%) than sites that produced fewer errors without URR ($n=$132;15%). We further analyzed instances where deploying URR produced more errors and made two notable observations. First, a majority of these errors refer to (a) rejections in setting cookies, (b) CSP violations, and (c) Timeout errors (potentially resulting from waiting for network responses). Second, 63% ($n=$84/132) of these instances involve a difference <=<=< =2 errors between the two visits. We note that our analysis is limited and is only one of many insights into page breakage. With that caveat, considering that a large majority of websites respond similarly with and without deploying URR (overall average, #sites with <=<=< = errors, #sites with $>=2$ errors), our observations show that URR retains web functionality and compatibility when deployed in the wild.
@@ 5. Discussion
Applicability to Other Systems
@@
In this work, we target and evaluate against JavaScript bundles generated by Webpack. We selected Webpack because it is the most popular bundling system on the Web. However, many other bundling approaches are used on the Web, some variations on the same approaches used in Webpack, some fundamentally different, and even working on different levels of the deployment process. We here briefly discuss these other bundling formats and how URR could be extended to apply to them (and with what difficulties).
@@
Webpack is generally used to combine multiple different JavaScript libraries and code units and to process them into a single JavaScript file to make deployment easier and (in some cases) execution faster. While Webpack is the most common tool for this purpose, many others are used with different languages, build chains, testing frameworks, and, in some cases, bundling additional resource types beyond JavaScript files. Examples of these alternative bundlers include Browserify  [^40] https://browserify.org/ and Gulp  [^41] https://gulpjs.com/, among many others.
@@
URR could easily be extended to cover these other bundling tools, as at root they all operate in the same manner (i.e., consume multiple JavaScript files, preprocess them, and then generate code for the resulting combined AST). To do so would only require generating new signatures for ASTs of each target privacy-harming library generated by the bundler’s preprocessing and rewriting phases and understanding the structure of each bundler combines the ASTs generated by each input library into the final, resulting code unit. This is work that would only need to be done once per bundler version and then could be shared across all URR clients.
@@
Another approach to JavaScript bundling is directly combining each JavaScript code unit into a single archive and shipping the entire archive to the client alongside the website’s initial HTML. This approach, exemplified by Google’s WebBundles  [^42] https://wpack-wg.github.io/bundled-responses/draft-ietf-wpack-bundled-responses.html proposal, does not preprocess or otherwise modify the include JavaScript code units; included files are directly copied into the bundle archive. Cloudflare’s Cloudflare’s Managed Components  [^43] https://managedcomponents.dev/ product can also be seen as a form of this kind of bundling, though instead of all combined into a single archive, they’re instead delivered “on demand”, as managed by an overriding “manager” application, making any URLs unpredictable. Extending URR to cover these kinds of bundled applications would be trivial since bundled JavaScript files are not modified in the bundling process. Identifying them within the bundled application is straightforward. Similarly, rewriting unwanted code is trivial since it only requires swapping the original file with the privacy-preserving alternative (either in the original archive or returned as a new subresource by the “manager” application).
Alternative Deployment Strategies
@@
In this work, we implement URR as a Firefox extension. We choose a Firefox extension because only Firefox’s extension API includes the ability for an extension to buffer and rewrite a fetched subresource  [^44] https://developer.mozilla.org/en-US/docs/Mozilla/Add-ons/WebExtensions/API/webRequest/filterResponseData (like a JavaScript file), before the file is seen, parsed, and executed by the JavaScript engine.
@@
However, the same approach we take in the extension could be identically deployed from other decision points, either to be more general across browsers (e.g., as part of a man-in-the-middle proxy) or more specific to particular browsers (e.g., as a modification to the V8 parsing pipeline). The majority of the resources needed for URR to work could apply equally across all of these intervention points (e.g., fingerprints of the ASTs of unwanted to code, bundle “unbundling” logic, privacy-preserving replacement AST subtrees).
@@
However, each of these intervention points would come with their own largely-predictable tradeoffs. Analyzing, unbundling, and rewriting JavaScript bundles as part of a man-in-the-middle proxy would work for any browser or network tool, though with a performance impact (since JavaScript engine optimizations like partial or lazy compilation would not be possible). Likewise, pushing URR logic into the JavaScript engine directly would likely allow for greater performance at the cost of development and maintenance cost, and require browser-specific implementations. That said, we reemphasize that most of the novel aspects of URR would be generic and shareable across all possible intervention points.
Diversity of Target Library Representation
@@
URR requires a significant amount of precomputation to work effectively. Specifically, URR requires precomputing a signature for every AST for each library or code unit that should be rewritten at runtime. While this is a significant improvement over the existing state of the art ((Smith et al., 2021), for example, requires the precomputing each target bundled application, which will both be orders of magnitude larger in occurrence, but also enormously larger in terms of required disk space), it is still not trivial. The same target library can give different signatures depending on library version (e.g., FingerprintJS v2 vs v3), bundler version (e.g., Webpack v4 vs Webpack v3), bundler optimization strategies (e.g., “tree-shaking” or no), library integration method (e.g.,CommonJS vs ECMAScript modules) among other dimensions.
@@
As discussed in Section 3, URR uses several heuristics to effectively generalize signatures, to flatten the number of dimensions of signatures needed per target library (e.g., label stripping, AST simplification, etc.). Nevertheless, there is still a tradeoff between coverage—generating as many signatures as possible to correctly identify the same target code across a wide range of representations—and concision—minimizing the memory, matching time, and disk space used on each URR client at runtime.
Replacements
While other phases of URR are programmatically generated, we manually created benign replacements. Even with manual curation, we show in Section 4.4.1 that while benign replacements can take between 20 and 120 minutes to create, they can be scaled to replace a large number of versions of a target library (e.g.,Sentry’s replacement applies to $>$300 versions).
@@
Regardless, this phase requires an expert understanding of both the specific bundler and the target library itself. While prior work has shown approaches that can automate this creation, we leave the adaption of a similar approach in the context of bundles as an avenue to explore in future work. Prior work like (Smith et al., 2021), which allows for the automatic creation of privacy-and-compatibility preserving versions of JavaScript libraries, could be leveraged to greatly expand the number of target libraries URR can identify and rewrite in bundled applications.
Caching Frequently Encountered Bundles
We designed URR to work against an extensive array of configuration options applied to bundles in the wild. Our deployment of URR makes an essential tradeoff: its large database of target library signatures consumes notable memory space (2.27 GB - 3.55 GB). In Section 4.3.5, we discussed that reducing the database size can affect URR’s detection of target libraries. However, developers can alternatively cache entire bundles that they frequently encounter. This way, while the browser extension would still block network requests, scripts do not need to be processed through URR on repeated encounters. For example, the service shopee.com has multiple related sites (shopee.com, shopee.co.id, shopee.co.th, shopee.tw), all of which load the same bundle (containing FPJS as one of the modules) from different URLs which would bypass filter lists. However, since the script’s content is identical, a cache of this script can help perform detection and replacement without needing to process the script into component modules.